OpenStudy (juliannamartineez):

Determine whether y varies with x. if so, find the constant of variation k and write an equation. x y 7 11 8 13 9 15 10 17

563blackghost (563blackghost):

\(\huge{y=kx}\) Plug them in and simplify if they do vary....they would each would need to have the same value of k.... \(\huge{11=k(7)}\) \(\huge{13=k(8)}\) \(\huge{15=k(9)}\) \(\huge{17=k(10)}\)

OpenStudy (juliannamartineez):

all values are going up by one..

563blackghost (563blackghost):

For the x-values, yes they are going up by 1. But we need to find k in each equation I typed. We would need to divide y by x to find k.... \(\huge{\frac{11}{7}=k}\) \(\huge{\frac{13}{8}=k}\) Once you get the totals from the both determine if they are the same...if they are not then `y does not vary directly to x`.

OpenStudy (juliannamartineez):

confused on how to divide them.. @563blackghost

OpenStudy (juliannamartineez):

divided 11 by 7?

563blackghost (563blackghost):

Correct :)

OpenStudy (juliannamartineez):

1.5

OpenStudy (juliannamartineez):

and 18.8 is 1.6

OpenStudy (juliannamartineez):

so y doesnt vary

OpenStudy (juliannamartineez):

@563blackghost